2. (30 pts) Use the method of undetermined coefficients to solve the following ODE y5y6y -12 e-2=, y(0) = 0, y(0) = -3
Answer #1

y +5y +by = -47 te Aurillary equation :D²+50+6=0 - D = -2,-3 ge=rey trze Чp = A + Ва е y= Be2Y_2B xe 2 4=-upé + lere 14 e